# Linzhan Mou5, Linqi Song4, Zhenan Sun [^2] Corresponding authors. 1,3, Ying Wei22footnotemark: 2 6 1

### v1

@@ Abstract
Quantizing large language models (LLMs) presents significant challenges, primarily due to outlier activations that compromise the efficiency of low-bit representation. Traditional approaches mainly focus on solving Normal Outliers—activations with consistently high magnitudes across all tokens. However, these techniques falter when dealing with Massive Outliers, which are significantly higher in value and often cause substantial performance losses during low-bit quantization. In this study, we propose DuQuant, an innovative quantization strategy employing rotation and permutation transformations to more effectively eliminate both types of outliers. Initially, DuQuant constructs rotation matrices informed by specific outlier dimensions, redistributing these outliers across adjacent channels within different rotation blocks. Subsequently, a zigzag permutation is applied to ensure a balanced distribution of outliers among blocks, minimizing block-wise variance. An additional rotation further enhances the smoothness of the activation landscape, thereby improving model performance. DuQuant streamlines the quantization process and demonstrates superior outlier management, achieving top-tier results in multiple tasks with various LLM architectures even under 4-bit weight-activation quantization. Our code is available at https://github.com/Hsu1023/DuQuant.
@@ 3.2 The proposed DuQuant Method
    Identify the feature dimension $d^{(1)}$ where the outlier are primarily concentrated, i.e., $d^{(1)}=\arg\max_{i}(\max_{j}|\mathbf{X}_{ij}|)$. Here, $\mathbf{X}_{ij}$ represents the element in the $i$-th row and $j$-th column of $\mathbf{X}$.

### v2

@@ Abstract
Quantization of large language models (LLMs) faces significant challenges, particularly due to the presence of outlier activations that impede efficient low-bit representation. Traditional approaches predominantly address Normal Outliers, which are activations across all tokens with relatively large magnitudes. However, these methods struggle with smoothing Massive Outliers that display significantly larger values, which leads to significant performance degradation in low-bit quantization. In this paper, we introduce DuQuant, a novel approach that utilizes rotation and permutation transformations to more effectively mitigate both massive and normal outliers. First, DuQuant starts by constructing rotation matrices, using specific outlier dimensions as prior knowledge, to redistribute outliers to adjacent channels by block-wise rotation. Second, We further employ a zigzag permutation to balance the distribution of outliers across blocks, thereby reducing block-wise variance. A subsequent rotation further smooths the activation landscape, enhancing model performance. DuQuant simplifies the quantization process and excels in managing outliers, outperforming the state-of-the-art baselines across various sizes and types of LLMs on multiple tasks, even with 4-bit weight-activation quantization. Our code is available at https://github.com/Hsu1023/DuQuant.
@@ 3.2 The proposed DuQuant Method
    Identify the feature dimension $d^{(1)}$ where the outlier are primarily concentrated, i.e., $d^{(1)}=\arg\max_{j}(\max_{i}|\mathbf{X}_{ij}|)$. Here, $\mathbf{X}_{ij}$ represents the element in the $i$-th row and $j$-th column of $\mathbf{X}$.